OpenStudy (anonymous):

Find the slope of a line that passes through (–2, –3), and (1, 1). A. 1/3 B. 1 C. 2 D. 4/3

OpenStudy (tkhunny):

\(\dfrac{Rise}{Run} = \dfrac{1 - (-3)}{1 - (-2)} = \dfrac{\Delta Y}{\Delta X}\)